# Unified Kraft Break at $\sim\,$6500 K: A Newly Identified Single-Star Obliquity Transition Matches the Classical Rotation Break

### v1

@@ 3 $T_{\rm eff}$  Boundary Determination
For each candidate cut, we required at least five systems in both the cool and hot populations to ensure adequate sample sizes. We then compared the $\lambda$ or $v\sin{i_{\star}}$  distributions of the two populations using the two-sample Kolmogorov-Smirnov (KS; Hodges 1958) test as implemented in scipy (Virtanen et al., 2020). We quantified the statistical significance as $\log_{10}(p)$, where $p$ is the KS test $p$-value, since the logarithmic scale facilitates comparison across small $p$-values. We defined the optimal boundary as the $T_{\rm eff}$  cut that maximized this quantity.
To estimate the uncertainty in the optimal boundary, we performed a bootstrap analysis. We generated 100,000100{,}000 bootstrap realizations by resampling the targets with replacement. In each realization, we perturbed every star’s $T_{\rm eff}$ by drawing from a Gaussian distribution centered on its reported value with standard deviation equal to its measurement uncertainty (i.e., $\sigma_{T_{\rm eff}}$). Note that a conservative 2.4% floor uncertainty in $T_{\rm eff}$ was adopted (Tayar et al., 2022). For each bootstrap sample, we repeated the boundary search over the same $T_{\rm eff}$ grid and identified the turning point that maximized the two-sample KS statistic. The resulting distribution of 100,000100{,}000 turning points was then used to derive the cut and the corresponding lower and upper uncertainties.
@@
4 Why the Break Looks Cooler in Binary Samples
@@
As noted in subsection 2.2, both theoretical models and observational evidence demonstrate that stellar companions can efficiently tilt planetary systems. Assuming a representative protoplanetary disk with a disk mass of $m_{\rm disk}=10^{-2}M_{\star}$, an inner edge at $a_{\rm in}=0.05\,\mathrm{AU}$ and an outer edge at $a_{\rm out}=50\,\mathrm{AU}$ following a surface density profile $\Sigma\propto r^{-1}$, together with an eccentricity of $e^{\prime}=0.5$ and a mutual inclination of $i^{\prime}=45^{\circ}$ between the companion’s orbit and the disk, we find [^6] The precession timescales are estimated by interpolating the contour scaling in Figure 3 of Batygin (2012), using representative points along the precession period curves. The code can be found here. that the disk precession timescales are less than, or comparable to, $10\,\mathrm{Myr}$ for all misaligned systems (except HAT-P-7, WASP-94, WASP-100, and WASP-180, which have wide-separation and low-mass stellar companions, see Table 1), implying that significant spin–orbit excitation by binary companions is possible (Batygin, 2012; Lai, 2014). This picture alone, however, does not explain the relatively sharp rise in the misaligned fraction at ${6105}\,\mathrm{K}$ (Figure 1, panel a) unless the companions’ properties (e.g., separations, mass ratios, eccentricities) change abruptly near this temperature. Current binary observations lack the temperature resolution needed to test for such a localized step and therefore cannot yet confirm this hypothesis (Offner et al., 2023).
@@
We observe that, although the single-star obliquity transition is determined to be $T_{\rm eff}={6447}$ K based on the optimal decision boundary for KS tests, there are two systems cooler than this threshold that are misaligned (Figure 1, panel b): WASP-101 ($T_{\rm eff}=6380\pm 120$ K, $\lambda=34^{\circ}\pm 3$; Zak et al. 2024), and XO-4 ($T_{\rm eff}=6397\pm 70$ K, $\lambda=46\fdg 7^{+6.1}_{-8.1}$; Narita et al. 2010). Our preceeding discussion suggests it to be plausible that these systems host as-yet-unidentified stellar companions. XO-4 has been observed with adaptive optics imaging (Adams et al., 2013) with no companion detected, whereas WASP-101 has not yet been imaged at high angular resolution. Interestingly, the measured $\lambda$ values for these two systems lie very close to $35^{\circ}$. If we assume $\lambda\approx\psi$, these values are consistent with one of two peaks ($\sim 35^{\circ}$ and $\sim 115^{\circ}$) in the obliquity distribution predicted from from von Zeipel-Kozai-Lidov cycle simulations (Fabrycky & Tremaine, 2007). This coincidence suggests that these systems may indeed host undetected stellar companions; and targeted follow-up observations will be required to confirm or refute this possibility.
@@
Either way, however, it remains puzzling how stellar companions could influence the rotation of the primary star, or tidal interactions with its planets, at the companion orbital separations relevant to our sample. Prior work shows that spin changes are most effective in close binaries through tidal interactions or companion-induced disk truncation/warping (i.e., at separations $\lesssim$ a few-tens of AU, Meibom et al. 2007). By contrast, most binaries in our sample have projected separations of order $10^{2}$-$10^{3}$ AU, in currently standard stellar-evolution recipes, systems at these separations are modeled as effectively isolated stars, with no meaningful cross-influence on interior structure or evolutionary tracks.
@@
Another possibility is that blended photometry/spectroscopy in binaries leads to systematic $T_{\rm eff}$ biases (e.g., Andersen 1991; Torres et al. 2010a; El-Badry et al. 2018; Furlan & Howell 2020; Koenigsberger et al. 2025), producing an apparent offset in the obliquity and rotation transitions relative to singles. Interestingly, we found that in our obliquity sample (consisting of hot Jupiters), the planet host is always the more massive component of a binary (See Figure 2). This is consistent with the higher hot-Jupiter occurrence around higher-mass stars (Johnson et al., 2010), but it may also reflect detection bias, since transiting planets are more readily found around the brighter primary (e.g., Wang et al. 2019). In such cases, it is tempting to suppose that a close, cooler companion with modest brightness contrast biases the composite spectrum/spectral energy distribution (SED) and drives the inferred $T_{\rm eff}$ downward, artificially placing these systems into the “cooler” regime. As shown in the Appendix A, however, the host stars’ $T_{\rm eff}$ values derived from multi-component SED modeling are consistent with the $T_{\rm eff}$ adopted in this work from TEPCat, indicating that contamination of the primary’s SED by a stellar companion cannot explain the $\sim 400~\mathrm{K}$ $T_{\rm eff}$ shift. It is worth noting, however, that the SED fitting relies on spectroscopically derived $\left[{\rm Fe}/{\rm H}\right]$ priors, which could still be biased by light contamination from stellar companions (El-Badry et al., 2018). Therefore, spatially resolved spectroscopic observations for the binaries listed in Table 1 are required to assess potential spectral contamination affecting the determination of the primaries’ $T_{\rm eff}$.
5 Implications of a higher Kraft break
5.1 Few RMs above the single-star Kraft break
@@
A central open question in the study of stellar obliquity, as recently reviewed by Albrecht et al. (2022), is whether spin-orbit misalignments are a byproduct of high-eccentricity migration, and therefore confined to hot-Jupiter-like systems (isolated close-in gas giants plausibly delivered by high-$e$ pathways), or instead reflect a universal process that is largely independent of planetary architecture.
@@
Although, for practical reasons, most RM measurements have targeted hot Jupiters, NASA’s K2 (Howell et al., 2014) and TESS (Ricker et al., 2015) missions have provided suitable targets to extend obliquity measurements beyond hot Jupiters and have revealed intriguing trends.
@@
The emerging picture has so far been that large spin-orbit misalignments concentrate in hot-Jupiter-like, isolated systems (including isolated sub-Saturns that may share high-$e$ delivery channels, Yu & Dai 2024), whereas other categories of planetary systems are typically aligned (see Figure 3), including compact multi-planet systems (Albrecht et al., 2013; Wang et al., 2018; Zhou et al., 2018; Wang et al., 2022; Dai et al., 2023; Lubin et al., 2023; Radzom et al., 2024, 2025), warm Jupiters (Wang et al., 2021; Rice et al., 2022; Wang et al., 2024), eccentric Jupiters (Espinoza-Retamal et al., 2023; Mireles et al., 2025), and very massive close-in companions with $m_{p}/M_{\star}\gtrsim 2\times 10^{-3}$ (about $2\,M_{\rm Jup}$ for a solar-mass star, Hébrard et al. 2011; Triaud 2018; Albrecht et al. 2022; Zhou et al. 2019; Rusznak et al. 2025), as well as brown dwarfs (e.g., Giacalone et al., 2024; Ferreira dos Santos et al., 2024; Brady et al., 2025; Carmichael et al., 2025; Doyle et al., 2025; Vowell et al., 2025; Zak et al., 2025). Taken at face value, this contrast may suggest that spin–orbit misalignment is not universally primordial across planetary architectures.
@@
Measuring the spin-orbit (mis)alignments of non-hot-Jupiter planets around cool stars, however, provides little discriminating power between these hypotheses, because hot Jupiters around cool stars are also predominantly aligned. A decisive test instead requires obliquity measurements to be taken of non-hot-Jupiter planets around hot hosts. Thus, shifting the single-star Kraft break to $\sim 6500\,\mathrm{K}$ — and, thereby, changing the boundary between which stars we classify as cool vs. hot — changes the inference we are permitted to draw from the measurements we have in hand. At present there are only five such systems (CoRoT-3, $T_{\rm eff}$$6558\pm 44$ K, $\lambda=-37\fdg 6^{+22.3}_{10.0}$, Triaud et al. 2009; TOI-3362, $T_{\rm eff}$$6800\pm 100$ K, $\lambda=1\fdg 2^{+2.8}_{-2.7}$ , Espinoza-Retamal et al. 2023; TOI-558, $T_{\rm eff}$$6466^{+95}_{-93}$ K, $\lambda=11\fdg 7^{+7.5}_{-4.8}$ , Espinoza-Retamal et al. 2025; TOI-778, $T_{\rm eff}$$6643\pm 150$ K, $\lambda=18\pm 11$ , Clark et al. 2023 and WASP-120, $T_{\rm eff}$$6450\pm 120$ K, $\lambda=-2\pm 4$ , Zak et al. 2024) with RM measurements around stars hotter than our single-star Kraft break, and only one of these (TOI-3362) lies beyond $1\sigma$ of the break (See Figure 3). With so few robust measurements around genuinely hot stars, the central question — as to whether spin-orbit misalignment is intrinsically more common around hotter stars, regardless of planet type — remains open.
@@
A further, albeit tentative, population-level line of evidence that hot stars may be primordially more misaligned comes from the distributions of $v\sin{i_{\star}}$ (Louden et al., 2021, 2024) and photometric modulation amplitudes (Mazeh et al., 2015): both are consistent with larger stellar tilts in hotter dwarfs, even for compact multi-planet systems. Moreover, among hot Jupiter hosts, misalignments appear more frequent around A-type than mid-F stars (Albrecht et al., 2022).
@@
One of the strongest remaining arguments against primordial spin-orbit misalignment comes from sub-Saturns around cool stars (see the bottom panel of Figure 3 ): single sub-Saturns around cool hosts are often misaligned (Winn et al., 2010b; Sanchis-Ojeda & Winn, 2011; Bourrier et al., 2022; Stefànsson et al., 2022; Knudstrup et al., 2024), whereas compact sub-Saturn multi-planet systems around similar cool hosts are typically well aligned (Radzom et al., 2024, 2025). This contrast indicates that the misalignments of single sub-Saturns around cool stars are most likely dynamical, not primordial. That conclusion, however, should not be naively extrapolated to hot Jupiters, because mechanisms that readily excite the obliquity of a low-mass sub-Saturn do not necessarily operate with comparable efficiency for a massive hot Jupiter.
@@
A concrete example of the perils of such extrapolation is secular resonance crossing (Petrovich et al., 2020), which can drive sub-Saturns onto polar orbits by transferring angular momentum deficit from an outer companion to an inner low-mass planet. Because the inner sub-Saturn carries less orbital angular momentum, its inclination can be strongly excited, whereas a hot Jupiter is much harder to tilt. This mechanism yields a testable prediction: around rapidly rotating hot stars, the stellar quadrupole can dominate over general-relativistic precession, detuning the resonance from exactly polar and driving a characteristic steady-state misalignment near $65^{\circ}$. Observationally, however, no sub-Saturn yet has an RM measurement above $\sim 6500\,\mathrm{K}$; recent attempts for TOI-1842 (Wittenmyer et al., 2022; Hixenbaugh et al., 2023) and TOI-1135 (Mallorquín et al., 2024; Dugan et al., 2025) target hosts that are hotter than the previously adopted $\sim 6100\,\mathrm{K}$ boundary but lie just below the single-star Kraft break at $\sim 6500\,\mathrm{K}$, leaving the hot-star prediction (sub-polar clustering near $\sim 65^{\circ}$) untested.
5.2 Obliquity Damping Mechanisms
@@
Our newly attained precision in determining the location of the obliquity transition, complementing that also achieved for the location of the Kraft break, may shed light on the physical origin of the obliquity transition itself. So far, two tidal mechanisms have been proposed for obliquity damping that predict an obliquity transition — the excitation of inertial waves which are then convectively dissipated (Lai, 2012; Saunders et al., 2024), vs. a net orbit-averaged tidal torque exerted by standing internal gravity waves (i.e., g-modes) which are resonance-locked with the orbital frequency (Zanazzi et al., 2024; Zanazzi & Chiang, 2025) over the course of the long-term evolution of the Brunt-Väisälä frequency. These proposals are associated with wave propagation and dissipation under different (although largely overlapping) regimes of stellar structure.
@@
Figure 4: Evolution of solar-composition MESA stellar models on either side of the convective-core transition. In both panels, the blue curves show the time evolution of $1.11~M_{\odot}$ stellar models, which do not exhibit a long-lived convective core on the main sequence, while the orange curves show those of $1.15~M_{\odot}$ stellar models, where a convective core persists throughout the main sequence until core hydrogen exhaustion. The upper panel shows the time evolution of the average Brunt-Väisälä frequency (or, equivalently, the g-mode undertone spacing), Equation 1, compared to a reference epoch of $t_{0}=1$ Gyr, as in Zanazzi et al. (2024). The lower panel shows how these stellar models evolve on the Hertzsprung-Russell diagram.
@@
We illustrate this feature of interest in the upper panel of Figure 4, using evolutionary stellar models on either side of this transition. These models [^7] Our MESA inlist files, and a shell script to generate the evolutionary tracks used in this work, have been made available as a deposit on https://doi.org/10.5281/zenodo.17470968 (catalog Zenodo). were generated with MESA r12778 (Paxton et al., 2011, 2013, 2015, 2018, 2019), using solar-calibrated values of the initial helium abundance and mixing length parameter. Both account for elemental diffusion, and include a small amount of convective overmixing, with $f_{\text{ov}}=0.05$. Over the course of main-sequence evolution, Zanazzi et al. (2024) and Zanazzi & Chiang (2025) show that the averaged Brunt-Väisälä frequencies of cool stars, and therefore all of their g-mode frequencies, increase significantly, gradually, and monotonically over the course of stellar evolution compared to their values $\left<N_{0}\right>=\left<N(t_{0})\right$> at some reference epoch $t_{0}$ of planet formation. By contrast, stars that are hotward of the Kraft break have g-modes whose frequencies actually decrease over substantial portions of their main-sequence lifetime, and rapidly increase near main-sequence turnoff. As a result, g-modes in these cool stars are capable of maintaining resonance locks (and, correspondingly, sustaining obliquity-damping torques) during the gradual inward migration of their planets, while those in stars hotward of the Kraft break are not. Thus, were resonance-locking tidal torques to be the dominant mechanism of obliquity damping, an obliquity transition should occur at the same temperature at which this transition in the behaviour of the g-mode frequencies also occurs.
@@
We now make the observation that this transition is distinct from the Kraft break. Both sets of evolutionary models shown in the lower panel of Figure 4 have very similar stellar masses ($1.11$ vs $1.15~M_{\odot}$), and very similarly sized convective envelopes, with quite significant depths (reaching to $0.76$ and $0.78~R_{\star}$, respectively). Rather, the cause of this transition is that the lower-mass, cooler, set of evolutionary models is not massive enough to possess a convective core, while the higher-mass set of evolutionary models sustains a convective core throughout its main-sequence lifetime. As in Zanazzi et al. (2024) and Zanazzi & Chiang (2025), our MESA modelling indicates that this transition occurs at a main-sequence effective temperature of roughly $6100\pm 100$ K, as can be seen in the lower panel of Figure 4. In general, the location of this stellar-mass boundary is also sensitive to other properties of the star, and to the physics adopted in stellar modelling, in particular the amount of overshooting/convective boundary mixing. However, it is known to be substantially cooler (being at a lower stellar mass) than the rotational Kraft break — e.g. Silva Aguirre et al. (2013); Bellinger et al. (2019); Buchele et al. (2025). So too, therefore, is the obliquity transition predicted by the resonance-locking mechanism alone.
@@
In practice, these two mechanisms are not likely to be mutually exclusive. Both may operate in stars which are cool enough to have convective envelopes and no convective cores ($T_{\rm eff}<6100$ K), and neither in stars with convective cores and no convective envelopes ($T_{\rm eff}>6500$ K). The two misaligned single-star systems discussed in Section 4, with $6100\!<\!T_{\rm eff}\!<\!6500$ K, then naturally fall into a regime where only inertial-wave tides operate; the damping is consequently weaker than in cooler stars where both channels are available. Nevertheless, the fact that the Kraft rotation break coincides with the single-star obliquity transition does at least strongly (if circumstantially) suggest that the loss of convective dissipation at higher stellar masses, rather than of resonance locking, is responsible for the obliquity transition.
6 Conclusion

### v2

@@ III $T_{\rm eff}$  Boundary Determination
For each candidate cut, we required at least five systems in both the cool and hot populations to ensure adequate sample sizes. We then compared the $\lambda$ or $v\sin{i_{\star}}$  distributions of the two populations using the two-sample Kolmogorov-Smirnov (KS; Hodges1958TheSP) test as implemented in scipy (virtanen2020scipy). We quantified the statistical significance as $\log_{10}(p)$, where $p$ is the KS test $p$-value, since the logarithmic scale facilitates comparison across small $p$-values. We defined the optimal boundary as the $T_{\rm eff}$  cut that maximized this quantity.
To estimate the uncertainty in the optimal boundary, we performed a bootstrap analysis. We generated 100,000100{,}000 bootstrap realizations by resampling the targets with replacement. In each realization, we perturbed every star’s $T_{\rm eff}$ by drawing from a Gaussian distribution centered on its reported value with standard deviation equal to its measurement uncertainty (i.e., $\sigma_{T_{\rm eff}}$). Note that a conservative 2.4% floor uncertainty in $T_{\rm eff}$ was adopted (Tayar2022). For each bootstrap sample, we repeated the boundary search over the same $T_{\rm eff}$ grid and identified the turning point that maximized the two-sample KS statistic. The resulting distribution of 100,000100{,}000 turning points was then used to derive the cut and the corresponding lower and upper uncertainties.
@@
IV Why the Break Looks Cooler in Binary Samples
@@
As noted in subsection II.2, both theoretical models and observational evidence demonstrate that stellar companions can efficiently tilt planetary systems. Assuming a representative protoplanetary disk with a disk mass of $m_{\rm disk}=10^{-2}M_{\star}$, an inner edge at $a_{\rm in}=0.05\,\mathrm{AU}$ and an outer edge at $a_{\rm out}=50\,\mathrm{AU}$ following a surface density profile $\Sigma\propto r^{-1}$, together with an eccentricity of $e^{\prime}=0.5$ and a mutual inclination of $i^{\prime}=45^{\circ}$ between the companion’s orbit and the disk, we find [^6] The precession timescales are estimated by interpolating the contour scaling in Figure 3 of Batygin2012, using representative points along the precession period curves. The code can be found here. that the disk precession timescales are less than, or comparable to, $10\,\mathrm{Myr}$ for all misaligned systems (except HAT-P-7, WASP-94, WASP-100, and WASP-180, which have wide-separation and low-mass stellar companions, see Table 1), implying that significant spin–orbit excitation by binary companions is possible (Batygin2012; Lai2014). This picture alone, however, does not explain why there is a relatively sharp change in the misaligned fraction at ${6105}\,\mathrm{K}$ (Figure 1, panel a).
@@
We observe that, although the single-star obliquity transition is determined to be $T_{\rm eff}={6447}$ K based on the optimal decision boundary for KS tests, there are two systems cooler than this threshold that are misaligned (Figure 1, panel b): WASP-101 ($T_{\rm eff}=6380\pm 120$ K, $\lambda=34^{\circ}\pm 3$; Zak2024A&A...686A.147Z), and XO-4 ($T_{\rm eff}=6397\pm 70$ K, $\lambda=46\fdg 7^{+6.1}_{-8.1}$; Narita2010). Our preceeding discussion suggests it to be plausible that these systems host as-yet-unidentified stellar companions. XO-4 has been observed with adaptive optics imaging (Adams2013) with no companion detected, whereas WASP-101 has not yet been imaged at high angular resolution. Interestingly, the measured $\lambda$ values for these two systems lie very close to $35^{\circ}$. If we assume $\lambda\approx\psi$, these values are consistent with one of two peaks ($\sim 35^{\circ}$ and $\sim 115^{\circ}$) in the obliquity distribution predicted from from von Zeipel-Kozai-Lidov cycle simulations (Fabrycky2007). This coincidence suggests that these systems may indeed host undetected stellar companions; and targeted follow-up observations will be required to confirm or refute this possibility.
@@
Either way, however, it remains puzzling how stellar companions could influence the rotation of the primary star, or tidal interactions with its planets, at the companion orbital separations relevant to our sample. Prior work shows that spin changes are most effective in close binaries through tidal interactions or companion-induced disk truncation/warping (i.e., at separations $\lesssim$ a few-tens of AU, Meibom2007). By contrast, most binaries in our sample have projected separations of order $10^{2}$-$10^{3}$ AU, in currently standard stellar-evolution recipes, systems at these separations are modeled as effectively isolated stars, with no meaningful cross-influence on interior structure or evolutionary tracks.
@@
Another possibility is that blended photometry/spectroscopy in binaries leads to systematic $T_{\rm eff}$ biases (e.g., Andersen1991; Torres2010; ElBadry2018; Furlan2020; Koenigsberger2025), producing an apparent offset in the obliquity and rotation transitions relative to singles. Interestingly, we found that in our obliquity sample (consisting of hot Jupiters), the planet host is always the more massive component of a binary (See Figure 2). This is consistent with the higher hot-Jupiter occurrence around higher-mass stars (Johnson2010), but it may also reflect detection bias, since transiting planets are more readily found around the brighter primary (e.g., Wang2019TESS). In such cases, it is tempting to suppose that a close, cooler companion with modest brightness contrast biases the composite spectrum/spectral energy distribution (SED) and drives the inferred $T_{\rm eff}$ downward, artificially placing these systems into the “cooler” regime. As shown in the Appendix A, however, the host stars’ $T_{\rm eff}$ values derived from multi-component SED modeling are consistent with the $T_{\rm eff}$ adopted in this work from TEPCat, indicating that contamination of the primary’s SED by a stellar companion cannot explain the $\sim 400~\mathrm{K}$ $T_{\rm eff}$ shift. It is worth noting, however, that the SED fitting relies on spectroscopically derived $\left[{\rm Fe}/{\rm H}\right]$ priors, which could still be biased by light contamination from stellar companions (ElBadry2018). Therefore, spatially resolved spectroscopic observations for the binaries listed in Table 1 are required to assess potential spectral contamination affecting the determination of the primaries’ $T_{\rm eff}$.
V Implications of a higher Kraft break
V.1 Few RMs above the single-star Kraft break
@@
A central open question in the study of stellar obliquity, as recently reviewed by Albrecht2022, is whether spin-orbit misalignments are a byproduct of high-eccentricity migration, and therefore confined to hot-Jupiter-like systems (isolated close-in gas giants plausibly delivered by high-$e$ pathways), or instead reflect a universal process that is largely independent of planetary architecture.
@@
Although, for practical reasons, most RM measurements have targeted hot Jupiters, NASA’s K2 (Howell2014) and TESS (Ricker2015) missions have provided suitable targets to extend obliquity measurements beyond hot Jupiters and have revealed intriguing trends.
@@
The emerging picture has so far been that large spin-orbit misalignments concentrate in hot-Jupiter-like, isolated systems (including isolated sub-Saturns that may share high-$e$ delivery channels, Yu2024), whereas other categories of planetary systems are typically aligned (see Figure 3), including compact multi-planet systems (Albrecht2013; Wang2018a; Zhou2018; WangX2022WASP148; Dai2023; Lubin2023; Radzom2024; Radzom2025), warm Jupiters (Wang2021; Rice2022WJs_Aligned; WangX2024), eccentric Jupiters (Espinoza2023; Mireles2025arXiv250922972M), and very massive close-in companions with $m_{p}/M_{\star}\gtrsim 2\times 10^{-3}$ (about $2\,M_{\rm Jup}$ for a solar-mass star, Hebrard2011; Triaud2018; Albrecht2022; Zhou2019; Rusznak2025), as well as brown dwarfs (e.g., Giacalone2024; Ferreira2024; Brady2025; Carmichael2025arXiv250618971C; Doyle2025; Vowell2025arXiv251000105V; Zak2025arXiv250520516Z). Taken at face value, this contrast may suggest that spin–orbit misalignment is not universally primordial across planetary architectures.
@@
Measuring the spin-orbit (mis)alignments of non-hot-Jupiter planets around cool stars, however, provides little discriminating power between these hypotheses, because hot Jupiters around cool stars are also predominantly aligned. A decisive test instead requires obliquity measurements to be taken of non-hot-Jupiter planets around hot hosts. Thus, shifting the single-star Kraft break to $\sim 6500\,\mathrm{K}$ — and, thereby, changing the boundary between which stars we classify as cool vs. hot — changes the inference we are permitted to draw from the measurements we have in hand. At present there are only five such systems (CoRoT-3, $T_{\rm eff}$$6558\pm 44$ K, $\lambda=-37\fdg 6^{+22.3}_{-10.0}$, Triaud2009; TOI-3362, $T_{\rm eff}$$6800\pm 100$ K, $\lambda=1\fdg 2^{+2.8}_{-2.7}$ , Espinoza2023; TOI-558, $T_{\rm eff}$$6466^{+95}_{-93}$ K, $\lambda=11\fdg 7^{+7.5}_{-4.8}$ , Espinoza2025; TOI-778, $T_{\rm eff}$$6643\pm 150$ K, $\lambda=18\pm 11$, Clark2023 and WASP-120, $T_{\rm eff}$$6450\pm 120$ K, $\lambda=-2\pm 4$, Zak2024A&A...686A.147Z) with RM measurements around stars hotter than our single-star Kraft break, and only one of these (TOI-3362) lies beyond $1\sigma$ of the break (See Figure 3). With so few robust measurements around genuinely hot stars, the central question — as to whether spin-orbit misalignment is intrinsically more common around hotter stars, regardless of planet type — remains open.
@@
A further, albeit tentative, population-level line of evidence that hot stars may be primordially more misaligned comes from the distributions of $v\sin{i_{\star}}$ (Louden2021; Louden2024) and photometric modulation amplitudes (Mazeh2015): both are consistent with larger stellar tilts in hotter dwarfs, even for compact multi-planet systems. Moreover, among hot Jupiter hosts, misalignments appear more frequent around A-type than mid-F stars (Albrecht2022).
@@
One of the strongest remaining arguments against primordial spin-orbit misalignment comes from sub-Saturns around cool stars (see the bottom panel of Figure 3 ): single sub-Saturns around cool hosts are often misaligned (Winnhatp11; Sanchis-Ojeda2011; Bourrier2022; Stefansson2022; Knudstrup2024), whereas compact sub-Saturn multi-planet systems around similar cool hosts are typically well aligned (Radzom2024; Radzom2025). This contrast indicates that the misalignments of single sub-Saturns around cool stars are most likely dynamical, not primordial. That conclusion, however, should not be naively extrapolated to hot Jupiters, because mechanisms that readily excite the obliquity of a low-mass sub-Saturn do not necessarily operate with comparable efficiency for a massive hot Jupiter.
@@
A concrete example of the perils of such extrapolation is secular resonance crossing (Petrovich2020), which can drive sub-Saturns onto polar orbits by transferring angular momentum deficit from an outer companion to an inner low-mass planet. Because the inner sub-Saturn carries less orbital angular momentum, its inclination can be strongly excited, whereas a hot Jupiter is much harder to tilt. This mechanism yields a testable prediction: around rapidly rotating hot stars, the stellar quadrupole can dominate over general-relativistic precession, detuning the resonance from exactly polar and driving a characteristic steady-state misalignment near $65^{\circ}$. Observationally, however, no sub-Saturn yet has an RM measurement above $\sim 6500\,\mathrm{K}$; recent attempts for TOI-1842 (Wittenmyer2022; Hixenbaugh2023) and TOI-1135 (Mallorquin2024; Dugan2025) target hosts that are hotter than the previously adopted $\sim 6100\,\mathrm{K}$ boundary but lie just below the single-star Kraft break at $\sim 6500\,\mathrm{K}$, leaving the hot-star prediction (sub-polar clustering near $\sim 65^{\circ}$) untested.
V.2 Obliquity Damping Mechanisms
@@
Our newly attained precision in determining the location of the obliquity transition, complementing that also achieved for the location of the Kraft break, may shed light on the physical origin of the obliquity transition itself. So far, two tidal mechanisms have been proposed for obliquity damping that predict an obliquity transition — the excitation of inertial waves which are then convectively dissipated (Lai2012; saunders_efficient_2024), vs. a net orbit-averaged tidal torque exerted by standing internal gravity waves (i.e., g-modes) which are resonance-locked with the orbital frequency (zanazzi2024damping; Zanazzi2025) over the course of the long-term evolution of the Brunt-Väisälä frequency. These proposals are associated with wave propagation and dissipation under different (although largely overlapping) regimes of stellar structure.
@@
Figure 4: Evolution of solar-composition MESA stellar models on either side of the convective-core transition. In both panels, the blue curves show the time evolution of $1.11~M_{\odot}$ stellar models, which do not exhibit a long-lived convective core on the main sequence, while the orange curves show those of $1.15~M_{\odot}$ stellar models, where a convective core persists throughout the main sequence until core hydrogen exhaustion. The upper panel shows the time evolution of the average Brunt-Väisälä frequency (or, equivalently, the g-mode undertone spacing), Equation 1, compared to a reference epoch of $t_{0}=1$ Gyr, as in zanazzi2024damping. The lower panel shows how these stellar models evolve on the Hertzsprung-Russell diagram.
@@
We illustrate this feature of interest in the upper panel of Figure 4, using evolutionary stellar models on either side of this transition. These models [^7] Our MESA inlist files, and a shell script to generate the evolutionary tracks used in this work, have been made available as a deposit on https://doi.org/10.5281/zenodo.17470968 (catalog Zenodo). were generated with MESA r12778 (mesa_paper_1; mesa_paper_2; mesa_paper_3; mesa_paper_4; mesa_paper_5), using solar-calibrated values of the initial helium abundance and mixing length parameter. Both account for elemental diffusion, and include a small amount of convective overmixing, with $f_{\text{ov}}=0.05$. Over the course of main-sequence evolution, zanazzi2024damping and Zanazzi2025 show that the averaged Brunt-Väisälä frequencies of cool stars, and therefore all of their g-mode frequencies, increase significantly, gradually, and monotonically over the course of stellar evolution compared to their values $\left<N_{0}\right>=\left<N(t_{0})\right$> at some reference epoch $t_{0}$ of planet formation. By contrast, stars that are hotward of the Kraft break have g-modes whose frequencies actually decrease over substantial portions of their main-sequence lifetime, and rapidly increase near main-sequence turnoff. As a result, g-modes in these cool stars are capable of maintaining resonance locks (and, correspondingly, sustaining obliquity-damping torques) during the gradual inward migration of their planets, while those in stars hotward of the Kraft break are not. Thus, were resonance-locking tidal torques to be the dominant mechanism of obliquity damping, an obliquity transition should occur at the same temperature at which this transition in the behaviour of the g-mode frequencies also occurs.
@@
We now make the observation that this transition is distinct from the Kraft break. Both sets of evolutionary models shown in the lower panel of Figure 4 have very similar stellar masses ($1.11$ vs $1.15~M_{\odot}$), and very similarly sized convective envelopes, with quite significant depths (reaching to $0.76$ and $0.78~R_{\star}$, respectively). Rather, the cause of this transition is that the lower-mass, cooler, set of evolutionary models is not massive enough to possess a convective core, while the higher-mass set of evolutionary models sustains a convective core throughout its main-sequence lifetime. As in zanazzi2024damping and Zanazzi2025, our MESA modelling indicates that this transition occurs at a main-sequence effective temperature of roughly $6100\pm 100$ K, as can be seen in the lower panel of Figure 4. In general, the location of this stellar-mass boundary is also sensitive to other properties of the star, and to the physics adopted in stellar modelling, in particular the amount of overshooting/convective boundary mixing. However, it is known to be substantially cooler (being at a lower stellar mass) than the rotational Kraft break — e.g. Silva2013; bellinger_inversion_2019; buchele_inversion_2025. So too, therefore, is the obliquity transition predicted by the resonance-locking mechanism alone.
@@
In practice, these two mechanisms are not likely to be mutually exclusive. Both may operate in stars which are cool enough to have convective envelopes and no convective cores ($T_{\rm eff}<6100$ K), whereas neither mechanism is expected to operate in stars with convective cores and no convective envelopes ($T_{\rm eff}>6500$ K). The two misaligned single-star systems discussed in Section 4, with $6100\!<\!T_{\rm eff}\!<\!6500$ K, then naturally fall into a regime where only inertial-wave tides operate; the damping is consequently weaker than in cooler stars where both channels are available. Nevertheless, the fact that the Kraft rotation break coincides with the single-star obliquity transition does at least strongly (if circumstantially) suggest that the loss of convective dissipation at higher stellar masses, rather than of resonance locking, is responsible for the obliquity transition.
VI Conclusion